# Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA

### v2

@@ Abstract
The power flow equations are central to many problems in power system planning, analysis, and control. However, their inherent non-linearity and non-convexity present substantial challenges during problem-solving processes, especially for optimization problems. Accordingly, linear approximations are commonly employed to streamline computations, although this can often entail compromises in accuracy and feasibility. This paper proposes an approach termed Conservative Bias Linear Approximations (CBLA) for addressing these limitations. By minimizing approximation errors across a specified operating range while incorporating conservativeness (over- or under-estimating quantities of interest), CBLA strikes a balance between accuracy and tractability by maintaining linear constraints. By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy. We illustrate the effectiveness of our proposed approach through several test cases, including its application to a unit commitment problem, where CBLA consistently achieves lower operating costs and improved feasibility compared to traditional linearization methods.
Index Terms:
Conservative bias linear approximation; power flow approximation; unit commitment.
@@ I Introduction
The power flow equations play a central role in the operation and analysis of electric power systems. These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7]. However, the nonlinearity of the power flow equations induces non-convexities in these problems that pose significant computational challenges.
To address these challenges, researchers have developed various linear approximations such as DC power flow [8], LinDistFlow [9], first-order Taylor expansions of the power flow equations, and other approximations [10]. These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems. However, these linearizations often depend on broad assumptions such as maintaining voltages near 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow. These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations. Consequently, the solutions derived from these linearized models may not closely align with the actual optimal solutions in real-world scenarios. This trade-off between simplicity and accuracy necessitates careful consideration when applying these linearizations in practice.
In response to these challenges, various studies have explored adaptive power flow approximations tailored to specific systems and operating ranges to enhance approximation accuracy (e.g., optimization-based approaches in [11] and sample-based approaches in [12, 13, 14]); see [15, 16, 17] for recent survey papers on this concept. For sample-based approaches, samples of operating points computed by repeatedly solving the power flow equations at various points within an operating range (e.g., a specified range of power injections) are leveraged to compute the approximations. By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest. These adaptive power flow linearizations spend computational effort to calculate the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems. By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
Extending the concept of sample-based adaptive power flow linearizations, the conservative linear approximation (CLA) approach in [12, 13] incorporates the concept of conservativeness. In other words, the CLAs are computed to minimize approximation errors with respect to the AC power flow equations while consistently over- or under-estimating quantities of interest over the set of drawn samples. The resulting approximations are particularly well suited for settings with an asymmetry in the implications of overestimating a quantity like voltage magnitude or current flow as opposed to underestimating that quantity. This is particularly relevant in power system optimization problems where feasibility is of paramount importance. For instance, when used in the bound on the magnitude of current flow through a line, a linearization that erroneously underestimates the amount of current flow risks predicting feasibility when the constraint is actually violated with respect to the nonlinear AC power flow equations. This is a more problematic linearization error than an overestimate of the current flow for use in this constraint. Thus, conservative linearizations that avoid errors in a particular direction (i.e., avoid either overestimates or underestimates of some quantity) are valuable in many power system optimization contexts. However, maintaining conservativeness can sometimes lead to reduced accuracy.
In this paper, we introduce an approach to approximating power flow equations called conservative bias linear approximation (CBLA). The CBLA approach seeks to balance the trade-off between conservativeness and accuracy, particularly in scenarios where certain samples are challenging to approximate accurately. To construct CBLAs, the process shares similarities with CLAs by beginning with drawing samples from within the operational range. These samples form the basis of a regression problem, which is solved to compute an approximated function representing the power flow equations. However, unlike CLA, CBLA does not explicitly enforce conservativeness in its approximated function as a hard constraint. Instead, the CBLA approach introduces an error function that penalizes linearization errors for samples that violate conservativeness to enable more accurate approximations.
CBLA offers the advantage of flexibility in designing customized error functions that quantify the penalty for deviating from actual values. User-defined error functions enable the approach to be tailored to particular quantities of interest and system characteristics, thus computing a linearization specialized for a specific problem. This flexibility can be particularly beneficial in scenarios where some violations are permissible, such as in chance-constrained optimization problems.
This work extends our previous study presented in [18], where we introduced the CBLA approach for power flow linearization. In this journal version, we build upon that foundation by incorporating a unit commitment formulation [19] using CBLA, comparing it against other power flow linearization methods, and detailing a step-by-step process for solving the unit commitment problem.
In summary, the paper’s main contributions are:
@@
    A CBLA formulation that is tailored to the specific system and operating range, optimal with respect to an error metric, and strikes a balance between conservativeness and accuracy.
@@
    A discussion on choosing an error function for computing the CBLA.
@@
    Numerical tests of CBLAs for a variety of test cases.
@@
    An application and numerical results on proposed unit commitment problems comparing the CBLA approach with other power flow linearizations.
The remainder of this paper is organized as follows: Section II presents background material on power flow equations and sample-based conservative linear approximations. Section III introduces the proposed conservative bias linear approximation (CBLA) approach. Section IV describes the iterative method for solving the unit commitment problem using CBLA. Section V provides numerical results demonstrating the performance of the proposed method. Finally, Section VI concludes the paper and outlines directions for future research.
II Background
In this section, we provide background information about the AC power flow equations and present the recently developed conservative linear approximations of these equations.
II-A The Power Flow Equations
Consider a power system where a reference bus has the voltage angle set to 00. Let $V(\theta)$ denote the voltage magnitude (angle). Let $P(Q)$ denote the active (reactive) power injection. We use the subscript $(\cdot)_{i}$ to represent a quantity at bus $i$ and the subscript $(\cdot)_{ik}$ to represent a quantity from or connecting bus $i$ to $k$. Let $j=\sqrt{-1}$. The AC power flow equations at bus $i$ are:
$P_{i}$ $=V_{i}^{2}G_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\cos\theta_{ik}+B_{ik}\sin\theta_{ik}),$ (1a)
$Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}),$ (1b)
where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G$ and $B$ are the real and imaginary parts of the system’s admittance matrix, respectively, and $\mathcal{B}_{i}$ is the set of all neighboring buses to bus $i$, including bus $i$.
@@ II-B Conservative linear approximations
The nonlinearity of the power flow equations in (1) contributes to the complexity encountered in solving optimization problems. To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig. 1)[12]. Moreover, CLAs facilitate parallel computation by enabling concurrent computation of the CLA for each quantity of interest. The construction of a CLA entails sampling power injections across an operational range of interest, such as a range of loads and power generated by Distributed Energy Resources (DERs), followed by computing power flow solutions for each sample and solving a constrained-regression problem.
For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$. This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }L_{d}\in\mathcal{N_{D}}\}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, $\mathcal{N_{D}}$ is the set of all buses with loads, and the superscripts max (min) indicate upper (lower) limits.
Figure 1: An illustration showcasing a comparison between a conventional linear approximation (on the left) and CLAs (on the right). In this visual representation, the solid line signifies the nonlinear function under consideration. In the figure on the left, the dotted line represents a traditional first-order Taylor approximation centered at point $\times$, while in the figure on the right, the dotted line above (below) corresponds to an over- (under-)estimating approximation.
The utilization of CLAs allows for the customization of the approximation to fit a specific operating range and the targeted system. Additionally, the sample-based approach enables integrating the behavior of complicated devices like tap-changing transformers and smart inverters into the approximation, as discussed in our prior work [6]. In the realm of optimization, CLAs offer a crucial advantage: they enable the satisfaction of nonlinear constraints while enforcing only linear inequalities, assuming the CLAs maintain conservativeness. Consequently, CLAs streamline optimization problems, rendering them suitable for commercial optimization solvers.
Consider a quantity of interest denoted as $\gamma$, which could represent variables such as the voltage magnitude at a specific bus or the magnitude of current flow along a particular line. In this context, bold quantities signify matrices and vectors. Let superscript $T$ denote the transpose. An overestimating CLA can be expressed as follows:
$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$ (2)
where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4). This CLA is designed to ensure the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
@@
Assuming that (3) is indeed satisfied, we can ensure that the constraint $\gamma\leq\gamma^{\text{max}}$ is also satisfied by instead enforcing a linear constraint $a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}\leq\gamma^{\text{max}}$. This approach allows us to meet the upper bound requirement $\gamma^{\text{max}}$ without introducing the implicit system of nonlinear AC power flow equations in (1). Importantly, by employing the CLA, we are able to satisfy the nonlinear equations while maintaining a linear formulation, thus enhancing computational tractability without sacrificing feasibility in the resulting solution.
To compute a CLA, we solve for the coefficients of the affine function of power injections in (2) in the following regression problem:
$\min_{a_{0},\ \bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}\mathcal{L}\left(\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right)\right)$ (4a)
@@
The subscript $(\,\cdot\,)_{m}$ denotes the $m^{\text{th}}$ sample and $M$ is the number of samples. The function $\mathcal{L}(\,\cdot\,)$ represents a loss function, such as the absolute value for $\ell_{1}$ loss or the square for squared-$\ell_{2}$ loss. In this paper, our focus is on quantities of interest denoted by $\gamma$, which correspond to the magnitudes of voltages ($V$) and current flows ($I$). The construction of underestimating CLAs follows a similar process as described in (4), with the key distinction being the reversal of the inequality direction in (4b).
The conservativeness of the CLA computed in (4) comes at the cost of reduced accuracy relative to the approximation corresponding to the unconstrained regression problem resulting from dropping (4a) from (4). To manage this tradeoff, the next section presents the main contribution of this paper, namely, a linear approximation technique that achieves a balance between conservativeness and accuracy. This approach involves biasing the linearization towards conservativeness, guided by a designated loss function.
@@ III Conservative Bias Linear Approximations
The CLA approach presented in Section II-B is consistently conservative within the set of drawn samples. However, in certain scenarios, the conservativeness property may lead to significant errors due to specific samples. In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range. The CBLA approach is designed to be optimal, minimizing a specific error metric while maintaining a tendency to be conservative to enhance accuracy. This means CBLA focuses on minimizing errors between the approximating function and the samples, allowing some samples to violate conservativeness at a specified cost.
@@ III-A Formulation
Let $\epsilon$ denote the mismatch between the approximated quantity and the actual quantity. The optimization problem to compute a CBLA is formulated as follows:
$\min_{f(\epsilon_{m}(a_{0},\bm{a}_{1}))}\quad\frac{1}{M}\sum_{m=1}^{M}f(\epsilon_{m})$ (5)
@@
$\epsilon_{m}=\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right),$ (6)
and
$f(\epsilon_{m})$ =$=\begin{dcases}g(\epsilon_{m}),&\text{if}\;\epsilon_{m}\leq 0\\ h(\epsilon_{m}),&\text{otherwise.}\\ \end{dcases}$ (7)
The optimization problem in (5) seeks to minimize the aggregated value of the error function $f(\,\cdot\,)$ defined in (7) over all samples by computing the coefficients $a_{0}$ and $\bm{a_{1}}$ in (6). This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$). The error function is computed based on the error’s sign for each sample. For overestimation, $h(\epsilon_{m})$ is set high to impose a substantial cost for violation, while $g(\epsilon_{m})$ is kept low for non-violations. Conversely, for underestimation, $h(\epsilon_{m})$ is relatively low and $g(\epsilon_{m})$ is high. For an alternative formulation of CBLA that incorporates a penalty term to influence the optimization behavior, see the appendix.
III-B Error function
Choosing a suitable error function in the CBLA approach is an important consideration. The choice of error function depends on various factors, such as the specific system requirements, a quantity of interest, and the trade-off between accuracy and conservativeness. To better understand how the error function works, we compare the error function used in the CLA approach with that of the CBLA approach.
The CLA approach imposes conservativeness across the set of sampled data in the constraints. We can rewrite the regression problem described in (4), which utilizes the $\ell_{1}$ loss function, as an optimization problem formulated in (5)–(7). In this formulation, the error function is defined as follows:
$f(\epsilon_{m})$ =$=\begin{dcases}\epsilon_{m},&\text{if}\;\epsilon_{m}\leq 0\\ \infty,&\text{otherwise.}\\ \end{dcases}$ (8)
Figure 2: An error function of the CLA where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
The error function in (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.

### v3

# Paprapee Buason12, Sidhant Misra1, Daniel K. Molzahn3

@@ Abstract
Accurate modeling of power flow behavior is essential for a wide range of power system applications, yet the nonlinear and nonconvex structure of the underlying equations often limits their direct use in large-scale optimization problems. As a result, linear models are frequently adopted to improve computational tractability, though these simplifications can introduce excessive approximation error or lead to constraint violations. This paper presents a linear approximation framework, referred to as Conservative Bias Linear Approximations (CBLA), that systematically incorporates conservativeness into the approximation process. Rather than solely minimizing local linearization error, CBLA constructs linear constraints that bound the nonlinear functions of interest over a defined operating region while reducing overall approximation bias. The proposed approach maintains the simplicity of linear formulations and allows the approximation to be shaped through user-defined loss functions tailored to specific system quantities. Numerical studies demonstrate that CBLA provides more reliable and accurate approximations than conventional linearization techniques, and its integration into a unit commitment formulation results in improved feasibility and reduced operating costs.
@@ I Introduction
Accurate representation of network physics is fundamental to a broad range of power system optimization problems. The nonlinear alternating current (AC) power flow equations govern voltage magnitudes, phase angles, and line flows, and therefore directly influence planning, operational scheduling, and reliability assessments. These equations appear in applications including resilient infrastructure design [gupta_buason_molzahn-fairness_pv_limit, owen_aquino_talkington_molzahn-EVacuation_feeder, bhusal2020], AC unit commitment (UC) [castillo2016, Sampath2018], and bilevel or market-based formulations [buason2023datadriven, wogrin2020]. However, embedding the full nonlinear equations into optimization models introduces nonconvexities that significantly complicate computation, particularly in large-scale or mixed-integer settings.
To mitigate these computational challenges, numerous linear approximations of the power flow equations have been developed. Widely used models such as DC power flow [stott2009], LinDistFlow [baran1989], and first-order Taylor expansions [wood2013power] provide tractable representations that enable efficient optimization. While these approximations are attractive due to their simplicity, they are typically derived under restrictive assumptions, for example, small voltage angle differences or near-nominal voltage magnitudes. When system operating conditions deviate from these assumptions, the resulting linear models may produce substantial errors, potentially leading to suboptimal or infeasible solutions when evaluated against the underlying AC equations. Thus, a fundamental trade-off arises between computational tractability and physical accuracy.
Recognizing these limitations, recent research has investigated adaptive linearization strategies that tailor approximations to a specific system and operating region. Rather than relying on globally valid but coarse assumptions, these methods construct linear models using information obtained from representative operating samples. Optimization-based techniques [misra2018optimal] and sample-driven approaches [BUASON2022, buason2024adaptive, CHEN2022108573] compute linear coefficients by minimizing deviations from AC power flow solutions within a prescribed range of power injections; see [jia2023tutorial1, 10202779, jia2023tutorial2] for surveys of this emerging direction. By shifting computational effort to an offline stage where samples are generated and processed, adaptive approximations can provide improved fidelity while preserving linear structure for subsequent optimization tasks.
Among these approaches, the conservative linear approximation (CLA) framework [BUASON2022, buason2024adaptive] incorporates an additional structural property: directional conservativeness. In CLA, the linear approximation is constructed so that, across the sampled operating region, quantities of interest are consistently overestimated or underestimated relative to the AC solution. This feature is particularly valuable in optimization problems where certain types of approximation error are more harmful than others. For example, underestimating line currents in a thermal limit constraint may falsely indicate feasibility, whereas overestimation results primarily in conservativeness. By enforcing one-sided error behavior, CLA reduces the risk of infeasible solutions. However, strict conservativeness may increase overall approximation error, especially in regions where the nonlinear relationship is difficult to represent accurately with a single linear model.
Motivated by this observation, this paper introduces a generalized framework termed conservative bias linear approximation (CBLA). The CBLA methodology retains the sample-based construction process but relaxes strict one-sided enforcement of conservativeness. Instead of imposing conservativeness as a hard constraint, CBLA incorporates it through a bias-aware error function that penalizes violations asymmetrically. This formulation enables a controlled balance between directional bias and approximation accuracy. As a result, CBLA can reduce excessive conservativeness while still discouraging the most critical forms of approximation error.
A distinguishing feature of CBLA is its flexibility in error function design. Users may define loss functions that reflect the operational priorities of a given application, allowing the approximation to emphasize particular quantities or error directions. This flexibility makes the framework adaptable to diverse optimization contexts, including those in which limited constraint violations are tolerable, such as chance-constrained or risk-aware formulations. By tailoring the regression objective to the needs of the downstream problem, CBLA produces linear models that are specialized to both the system and the intended optimization task.
This work extends our preliminary study in [buason_cbla], where the CBLA concept was first introduced for power flow approximation. In the present journal version, we expand the framework and demonstrate its practical implications through extensive numerical evaluation. In particular, we integrate CBLA into a UC formulation [Padhy2004], compare its performance with alternative linearization techniques, and describe a structured solution procedure for solving the resulting problem. The results illustrate that CBLA achieves improved feasibility and reduced operating cost relative to conventional linear models.
The primary contributions of this paper are summarized as follows:
@@
    A CBLA framework that balances directional conservativeness and approximation accuracy over a specified operating range.
@@
    A discussion of error function selection and its role in shaping the trade-off between bias and fidelity.
@@
    Numerical validation of CBLA across multiple test systems.
@@
    An application of CBLA to UC, including comparative results against established linear power flow models.
The remainder of this paper is organized as follows. Section II reviews relevant background on power flow modeling and conservative linear approximations. Section III presents the proposed CBLA methodology. Section IV details the UC formulation and solution approach. Section V provides numerical results, and Section VI concludes the paper.
II Background on AC Power Flow and Conservative Linear Approximations
This section reviews the AC power flow model and summarizes the sample-based conservative linear approximation (CLA) framework that motivates the development of the proposed approach.
II-A AC Power Flow Model
We consider a power network in which one bus is selected as the reference with voltage angle fixed at zero. Let $V_{i}$ and $\theta_{i}$ denote the voltage magnitude and phase angle at bus $i$, respectively. The active and reactive power injections are represented by $P_{i}$ and $Q_{i}$. Subscripts $(\cdot)_{i}$ refer to quantities associated with bus $i$, while $(\cdot)_{ik}$ indicates quantities related to the branch connecting buses $i$ and $k$.
The steady-state AC power flow equations at bus $i$ are given by
$P_{i}$ $=V_{i}^{2}G_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}\left(G_{ik}\cos\theta_{ik}+B_{ik}\sin\theta_{ik}\right),$ (1a)
$Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}\left(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}\right),$ (1b)
where $\theta_{ik}:=\theta_{i}-\theta_{k}$, and $G$ and $B$ denote the real and imaginary parts of the network admittance matrix. The set $\mathcal{B}_{i}$ contains bus $i$ and all buses directly connected to it.
The nonlinear trigonometric and multiplicative terms in (1) introduce nonconvexities that significantly complicate optimization problems embedding AC network constraints. This motivates the construction of tractable linear representations of selected quantities derived from these equations.
@@ II-B Conservative Linear Approximations
To improve tractability while preserving feasibility with respect to the AC model, we previously developed a sample-based CLA framework [BUASON2022]. The goal of CLA is to construct affine functions of power injections that approximate nonlinear quantities of interest—such as voltage magnitudes or current flows—while ensuring one-sided error behavior over a prescribed operating region.
The CLA procedure begins by defining an operating set $\mathcal{S}$ for power injections. For example, load demands may be sampled from a probability distribution $\mathbb{P}_{\mathcal{S}}$ over
$\mathcal{S}=\{P_{L_{d}}^{\min}\leq P_{L_{d}}\leq P_{L_{d}}^{\max},\;Q_{L_{d}}^{\min}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\max},\;\forall L_{d}\in\mathcal{N}_{D}\},$
where $\mathcal{N}_{D}$ denotes buses with loads and the superscripts indicate lower and upper bounds. A uniform distribution is commonly adopted. For each sampled injection vector, the AC power flow equations are solved to obtain the corresponding nonlinear quantities of interest. These samples are then used to compute an affine approximation through constrained regression.
An illustrative comparison between a conventional first-order Taylor expansion and a CLA is shown in Fig. 1. Unlike a local linearization centered at a single operating point, a CLA is constructed to systematically overestimate or underestimate the nonlinear function throughout the sampled region.
Figure 1: Comparison of a conventional linearization (left) and CLAs (right). The solid curve represents the nonlinear function. The dotted line on the left corresponds to a first-order Taylor approximation at the marked point. The dotted lines on the right represent over- and under-estimating conservative approximations.
In addition to improved regional fidelity, CLAs provide two practical advantages. First, each quantity of interest can be approximated independently, enabling parallel computation of coefficients. Second, if conservativeness is maintained, nonlinear inequality constraints can be replaced with linear inequalities while preserving feasibility relative to the AC model. This enables compatibility with commercial linear and mixed-integer solvers.
Let $\gamma$ denote a scalar quantity derived from the AC power flow solution, such as a voltage magnitude or branch current magnitude. Using vector notation for power injections, an overestimating CLA takes the affine form
$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$, (2)
where $a_{0}$ and $\bm{a}_{1}$ are regression coefficients. Conservativeness requires that, for injections within the operating range,
@@
If (3) holds, then an upper bound constraint $\gamma\leq\gamma^{\max}$ can be enforced through the linear condition
$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}\leq\gamma^{\max},$
without explicitly embedding the nonlinear AC equations. Thus, feasibility with respect to the nonlinear model is preserved while maintaining linear structure.
The coefficients in (2) are obtained by solving the constrained regression problem
@@
s.t. $\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right)\leq 0,\quad m=1,\dots,M.$ (4b)
where $m$ indexes samples, $M$ denotes the number of samples, and $\mathcal{L}(\cdot)$ is a loss function such as the absolute value ($\ell_{1}$) or squared error ($\ell_{2}$). Underestimating CLAs are obtained by reversing the inequality in (4b).
While the constraint in (4b) guarantees directional conservativeness, it restricts the feasible regression set and may increase approximation error compared to the unconstrained formulation obtained by removing this condition. Consequently, strict conservativeness can reduce overall accuracy, particularly in regions where the nonlinear relationship is difficult to capture with a single affine model.
The next section introduces the main contribution of this paper, which extends the CLA framework by incorporating directional bias directly into the regression objective. This formulation enables a more flexible trade-off between feasibility protection and approximation accuracy.
@@ III Conservative Bias Linear Approximations
The CLA framework described in Section II-B guarantees strict one-sided conservativeness across all sampled operating points. While this property is desirable for feasibility protection, enforcing it uniformly may lead to unnecessarily large approximation errors when certain samples are difficult to represent accurately with a single affine model.
To address this limitation, we introduce a sample-based conservative bias linear approximation (CBLA) methodology. Similar to CLA, the proposed approach is adaptive in that it is constructed using samples drawn from a specified operating region for a particular system. However, instead of imposing conservativeness as a hard constraint, CBLA incorporates a bias toward conservativeness directly into the optimization objective. This formulation enables the approximation to remain predominantly conservative while allowing limited violations at a controlled cost. As a result, CBLA seeks to achieve a principled balance between directional bias and approximation accuracy.
@@ III-A Formulation
Let $\epsilon_{m}$ denote the approximation error (mismatch) at sample $m$, defined as the difference between the true quantity of interest and its affine approximation. The CBLA coefficients are computed by solving the optimization problem
$\min_{a_{0},\,\bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}f(\epsilon_{m}),$ (5)
@@
[…]
The function $f(\epsilon_{m})$ determines how errors are penalized and is defined piecewise as
$f(\epsilon_{m})=\begin{dcases}g(\epsilon_{m}),&\text{if }\epsilon_{m}\leq 0,\\ h(\epsilon_{m}),&\text{otherwise.}\end{dcases}$ (7)
The optimization problem in (5) therefore minimizes the average penalized error across all samples by selecting coefficients $a_{0}$ and $\bm{a}_{1}$. The key distinction from CLA lies in the structure of the error function. Rather than forbidding violations of conservativeness, CBLA assigns asymmetric penalties based on the sign of $\epsilon_{m}$.
For an overestimating approximation, violations correspond to $\epsilon_{m}>0$. In this case, $h(\epsilon_{m})$ is chosen to impose a relatively large penalty, while $g(\epsilon_{m})$ remains comparatively small for non-violating samples. For underestimating approximations, the roles of $g(\cdot)$ and $h(\cdot)$ are reversed. This asymmetric design induces a tendency toward conservativeness without strictly enforcing it.
III-B Error Function Design
The selection of the error function $f(\cdot)$ plays a central role in determining the trade-off between conservativeness and approximation accuracy. Its design may depend on system requirements, the specific quantity being approximated, and the acceptable level of directional violation.
To illustrate the relationship between CLA and CBLA, consider the CLA regression problem in (4) with an $\ell_{1}$ loss. That formulation can be interpreted within the CBLA framework by defining
$f(\epsilon_{m})=\begin{dcases}\epsilon_{m},&\text{if }\epsilon_{m}\leq 0,\\ \infty,&\text{otherwise.}\end{dcases}$ (8)
Figure 2: Error function corresponding to CLA, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
As shown in Fig. 2, assigning infinite cost to $\epsilon_{m}>0$ strictly prohibits violations, thereby recovering the CLA formulation. In contrast, CBLA replaces the infinite penalty with a finite but potentially large function, permitting limited violations when doing so improves overall accuracy.